# Gravitational wave probes of particle dark matter: a review

### v4

@@ VI GW constraints on boson clouds
The extensive theoretical background on boson clouds around rotating BHs, coupled with the potential for them to be DM, have motivated GW probes of boson cloud/BH systems. Different GW interferometers probe vastly different mass regimes of ultralight dark matter boson clouds. While ground-based detectors are sensitive to annihilation signals from bosons of masses between $\sim 10^{-14}-10^{-12}$ eV Arvanitaki et al. (2015a), space-based detectors can probe a few orders of magnitude lower than that, $\sim 10^{-20}-10^{-16}$ eV Baumann et al. (2019a), and HFGW detectors could see systems with masses around $\sim 10^{-9}$ eV Aggarwal et al. (2021). At the moment, only ground-based GW interferometers exist, and so we focus on methods and search results from the most recent LIGO, Virgo and KAGRA observing runs. In particular, only GWs from annihilating boson clouds around isolated BHs, discussed in Section V.1, can be searched for and constrained with current GW data.
@@ VI.1 Search targets
    Known galactic BHs (e.g. Cygnus X-1 Fabian et al. (1989))
@@
    Known remnants of binary black hole mergers (e.g. GW150914 Abbott et al. (2016))
@@
    Known binary black hole inspirals and mergers Abac et al. (2025b)
@@
Each of these sources has its own benefits and drawbacks: known BHs have many of their intrinsic parameters measured – mass, spin, inclination angle, etc. –; however, they are old and thus the strength of GW emission from annihilating boson clouds would be significantly diminished with respect to that when they were first born. On the contrary, remnants of BH mergers are extremely young: their clouds build up quickly and they could begin emitting days or weeks after formation, as described in Section V. Their age is thus precisely measured, as well as the final mass and spin (within $10\%$ in the case of GW150914 Abbott et al. (2016)). These estimates can be even more precise than those that arise from some x-ray binaries. However, remnants of binary black hole mergers are far: GW150914 merged at $0.410$ Mpc, and other detected systems have been seen out to Gpcs. Thus, the scaling of GW amplitude with distance implies that we could not detect remnants of mergers with annihilating scalar boson clouds around them. We note, however, that rapidly annihilating vector boson clouds could be detected out to Gpc Jones et al. (2023, 2025) in current-generation GW interferometers, making merger remnants a viable target even now.
On the other hand, with known binary black hole inspirals, boson clouds could alter the GW signal in a variety of ways, as detailed in Section V.5. Thus, using the spins and masses inferred from LIGO, Virgo and KAGRA measurements, the allowed boson mass can be constrained: if certain BH masses are detected with low spins, this could be evidence that boson clouds have spun down BHs. On the contrary, if BHs with high spins are detected, this could rule out the presence of boson clouds existing in a particular mass range Ng et al. (2021a). Such an analysis allows us to obtain some evidence for or rule out the presence of boson clouds around rotating binary black holes, but does not necessarily permit us to claim confidently that we have detected a boson cloud.
Finally, unknown galactic BHs could take on a wide range of masses, ages and spins depending on when and how they formed. The number of expected BHs in the galaxy is $\mathcal{O}(10^{9})$, which comes from considering the age of the galaxy ($\sim 10^{11}$ years) and the supernova rate ($\sim 1$ per century) Madau and Dickinson (2014). Thus, the BHs we observe are such a small portion of the total number that exist in the local universe, and may not be representative of the full population. However, searches for such systems are entirely blind: we do not know any BH mass or spin parameters, nor do we know the location of such systems. While this may seem to be a daunting task, computationally efficient methods from the CW community have been leveraged to do these kinds of searches D’Antonio et al. (2018); Isi et al. (2019); Jones et al. (2023, 2025).
@@ VI.2 Search methods
VI.2.1 Excess power method
One approach to searching for boson clouds around rotating BHs is the excess-power method, which models the expected signal as approximately monochromatic over the entire observation time. Because the source could lie anywhere on the sky, the signal is modulated by the Earth’s motion, leading to a Doppler-shifted frequency. To recover such a signal, the data are transformed into time–frequency maps and corrected for different sky positions, so that a truly monochromatic signal would appear as a narrow line D’Antonio et al. (2018). This strategy is similar to that described in Section III.1.2, but here the modulation arises from the Doppler effect rather than from the random velocity distribution of individual DM particles. In practice, the analysis is repeated using different coherence times $T_{\text{FFT}}$, which control the width of the frequency bins. Shorter $T_{\text{FFT}}$ values are less sensitive overall but more robust to small frequency drifts, making the search less dependent on specific signal-model assumptions.
@@ VI.2.2 Viterbi
Another method that is more robust against stochastic frequency variations relies on Hidden Markov Models (HMMs) to find tracks in the time-frequency plane. The main idea is to model a time-frequency GW signal probabilistically as a Markov chain of transitions between “hidden” (unobservable) frequency states and use a detection statistic to relate “observed” frequency states with the hidden ones Quinn and Hannan (2001). In the framework for this method, the probability to jump to a new state at a given time only depends upon the previous hidden state. In total, there would be $N_{Q}^{N_{T}+1}$ possible paths through the hidden states, where $N_{Q}$ is the total number of hidden states and $N_{T}$ is the total number of times at which we have observed frequency states. The Viterbi algorithm provides a recursive, computationally efficient way of maximizing the probability that the hidden set of states is responsible for the observed sequence of states.
In Fig. 16, taken from Isi et al. (2019), we show the optimal path found by the Viterbi method for injected boson cloud signals in Gaussian noise that has an associated random walk of the frequency, which is meant to simulate unknown theoretical features of the signal. We can see that in both cases, with less and more variability of the signal, the Viterbi algorithm can find the appropriate track.
Recently, the Viterbi algorithm has also been adapted to search for GWs from annihilating vector boson clouds, which have timescales significantly shorter than in the scalar case Jones et al. (2023, 2025), and for which numerical waveforms must be used to gauge sensitivity Siemonsen and East (2020); Siemonsen et al. (2023b).
@@
Figure 16: Taken from Isi et al. (2019). This figure illustrates how the Viterbi HMM algorithm described in Section VI.2.2 works to identify tracks in the time-frequency plane that would arise from GW emission from annihilating boson clouds around rotating BHs. The physics behind this form of DM has been discussed in Section V.1. In particular, these plots show two examples of injected and recovered boson cloud signals with random walk frequency variations $|\delta f|\leq 0.1\Delta f$ (weak variation) and (b) $|\delta f|\leq 0.5\Delta f$ (strong variation), where $\Delta f=1/T_{\rm drift}$ is the frequency bin size for the coherence time of the analysis $T_{\rm drift}=8$ days. The top panels show the injected random walk frequency evolutions, which cannot be seen by eye in the time-frequency tracks. The Viterbi optimal path results in a good match to the injected signal path. Each step in time on the $x$-axis is 8 days. The signal has a frequency of $201.2$ Hz, a spin-up of $10^{-12}$ Hz/s, a randomly chosen sky position, and a strain amplitude of $5\times 10^{-26}$.
@@ VI.2.3 Cross-correlation and other methods
We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition of cosmological or astrophysical GW signals. Furthermore, virtually any method in the CW community Riles (2023) could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
@@ VI.3.1 All-sky search constraints
The excess power method described in Section VI.2.1 has been used to analyze LIGO O3 data to search for the presence of annihilating scalar boson clouds around isolated BHs within our galaxy Abbott et al. (2022d, e). This search was very computationally expensive, since each sky position had to be analyzed individually. No evidence for scalar boson cloud systems was found in either case, and thus upper limits were placed on the presence of boson clouds with different ages, spins, and distances away from earth in two different ways Palomba et al. (2019); Dergachev and Papa (2019); Abbott et al. (2022d):
@@
Figure 17: Taken from Abbott et al. (2022d). Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, exclusion region of particular boson mass /BH mass pairs are derived assuming different ages and distances (left: 1kpc; right: 15 kpc) for BHs with initial spins of $\chi_{i}=0.9$. The physics behind this form of DM has been discussed in Section V.1. This is the most model-agnostic way to probe boson clouds around rotating BHs: the limits do not assume distributions over the mass, spin, distance of BHs in the Galaxy.
Constraints from the second way are shown in Fig. 18, in which a Kroupa BH mass distribution $f(m)\propto M_{\rm BH}^{-2.3}$ Elbert et al. (2018) between $5,~50]M_{\odot}$ and $5,~100]M_{\odot}$ is assumed, as well as a uniform initial BH spin distribution between [0.2,0.9]. For a fixed BH age, the maximum distance reach as a function of boson mass indicates that at least 5% of GWs from BH/boson cloud systems would have been detected by the search, i.e. their strain amplitudes are larger than the minimum detectable amplitudes at a given confidence level (the upper limits) of the search (an example of this quantity is in Section III.1.2, Eq. 71. ). As expected, on average when $50M_{\odot}$ is the maximum BH mass considered, the distance that can be reached is also smaller, since the GW amplitudes scales strongly with the BH mass.
Figure 18: Taken from Abbott et al. (2022d). Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, the maximum distance that the search in Abbott et al. (2022d) could reach at which at least 5% of a simulated BH population would have been detectable, that is, with a strain greater than the value of the upper limit, is shown here. The physics behind this form of DM has been discussed in Section V.1. To create this plot, Ref. Abbott et al. (2022d) assumed a uniform spin distribution between [0.2,0.9], different ages, and a Kroupa mass distribution with maximum mass $50M_{\odot}$ (left) and $100M_{\odot}$ (right). These limits depend on the distributions assumed and are thus less agnostic than those presented in Fig. 17.
We note that the search constraints present in, for example, Steltner et al. (2023); Ming et al. (2024); Covas et al. (2025), could also be cast in terms of all-sky or directed search constraints on scalar or vector boson clouds, but with particular caveats that would have be worked out depending on the actual search parameter space.
@@ VI.3.2 Cygnus X-1 constraints
Cygnus X-1 is an x-ray binary that is close by, has a large spin, $\sim 0.9$, and has relatively well-measured orbital parameters, which make it a reasonable BH to target. Since the sky position is known, searches based on the Viterbi algorithm discussed in Section VI.2.2 can use much longer coherence times than those used in Abbott et al. (2022d), thus permitting exquisite sensitivity towards almost monochromatic signals, while also being robust against theoretical uncertainties in the GW emission from Cygnus X-1, e.g. due to uncertainties on the measured semi-major axis, etc., due to the presence of an accretion disk, etc.
@@
Constraints on self-interacting DM using the search results on O2 data in Sun et al. (2020) are shown only in the left-hand plot, since that search was designed to analyze 211 annihilation from Cygnus X-1. Furthermore, current and future ground-based interferometers will be able to place constraints on the axion decay constant at frequencies of $\mathcal{O}(100)$ Hz (left plot), while space-based interferometers, such as DECIGO and MAGIS-Space are required to probe transition signals at frequencies below 1 Hz (right panel). The figure also highlights the possibility of multi-band GW astronomy: it would be possible to detect a transition from the 322 to 211 energy levels in space-based observatories while simultaneously observing the CW arising from annihilation of bosons in the 211 energy level in ground-based interferometers Collaviti et al. (2025). Annihilation signals from the 322 level (middle plot) would be too weak to be detected by any of the considered current or future GW interferometers.
@@
Figure 19: Taken from Sun et al. (2020). Upper limits at 95% confidence on the strain amplitude of annihilating scalar boson clouds coming from the BH in Cygnus X-1. The physics behind this form of DM has been discussed in Section V.1. The black curve indicates the upper limits from the search on O2 data, with the red stars denoting upper limits derived in frequency bands with injections. The two different curves correspond to the signal amplitude with different choices for the (unknown) age of the system (lower curve: $10^{6}$ years; upper curve: $5\times 10^{6}$ years. The GW fine structure constant is colored, and the gray regions denote boson masses disfavored under the two different age assumptions. The parameters of Cygnux X-1 used to compute the amplitude curves are: $M=14.8M_{\odot}$, $\chi_{i}=0.99$, and $d=1.86$ kpc.
@@
Figure 20: Taken from Collaviti et al. (2025). Parameter space reachable by current and next-generation GW interferometers for (left) 211 annihilation, (middle) 322 annihilation, and (right) 322 $\rightarrow$ 211 transition signals from a scalar boson cloud around Cygnus X-1, assuming an age of $10^{5}$ yr and the parameters for Cygnus X-1 given in Fig. 19. The numbers 211, 322, and the 322 $\rightarrow$ 211 transition refer to specific energy levels of the cloud. The physics behind this form of DM has been discussed in Sections V.1 and V.2. The dashed curves divide the parameter space into three regimes: (1) top, above orange-dashed line: the strong self-interaction regime, (2) middle, between dashed-orange and dashed-blue lines: the moderate self-interaction regime, and (3) bottom, below the dashed-blue line: the gravitational (negligible self-interactions) regime. The white dotted line shows the expected relation between the axion and its decay constant. Darker colors here correspond to weaker GW signals primarily due to smaller occupation numbers in the energy states of the boson clouds. Constraints on 211 annihilation come from ground-based interferometers, while constraints from transitions will come from space-based interferometers. The 322 annihilation signal is too weak to be detected by any current our future observatory. The O2 constraints come from a real analysis of GW data and are thus representative of both the actual power spectral density of that dataset and an exclusion region on the plot.

### v5

@@ VI GW constraints on boson clouds
The extensive theoretical background on boson clouds around rotating BHs, coupled with the potential for them to be DM, have motivated GW probes of boson cloud/BH systems. Different GW interferometers probe vastly different mass regimes of ultralight dark matter boson clouds. While ground-based detectors are sensitive to annihilation signals from bosons of masses between $\sim 10^{-14}-10^{-12}$ eV Arvanitaki:2014wva, space-based detectors can probe a few orders of magnitude lower than that, $\sim 10^{-20}-10^{-16}$ eV Baumann:2018vus, and HFGW detectors could see systems with masses around $\sim 10^{-9}$ eV Aggarwal:2020olq. At the moment, only ground-based GW interferometers exist, and so we focus on methods and search results from the most recent LIGO, Virgo and KAGRA observing runs. In particular, only GWs from annihilating boson clouds around isolated BHs, discussed in Section V.1, can be searched for and constrained with current GW data.
@@ VI.1 Search targets
    Known galactic BHs (e.g. Cygnus X-1 Fabian:1989ej)
@@
    Known remnants of binary black hole mergers (e.g. GW150914 Abbott:2016blz)
@@
    Known binary black hole inspirals and mergers LIGOScientific:2025hdt
@@
Each of these sources has its own benefits and drawbacks: known BHs have many of their intrinsic parameters measured – mass, spin, inclination angle, etc. –; however, they are old and thus the strength of GW emission from annihilating boson clouds would be significantly diminished with respect to that when they were first born. On the contrary, remnants of BH mergers are extremely young: their clouds build up quickly and they could begin emitting days or weeks after formation, as described in Section V. Their age is thus precisely measured, as well as the final mass and spin (within $10\%$ in the case of GW150914 Abbott:2016blz). These estimates can be even more precise than those that arise from some x-ray binaries. However, remnants of binary black hole mergers are far: GW150914 merged at $0.410$ Mpc, and other detected systems have been seen out to Gpcs. Thus, the scaling of GW amplitude with distance implies that we could not detect remnants of mergers with annihilating scalar boson clouds around them. We note, however, that rapidly annihilating vector boson clouds could be detected out to Gpc Jones:2023fzz; Jones:2024fpg in current-generation GW interferometers, making merger remnants a viable target even now.
On the other hand, with known binary black hole inspirals, boson clouds could alter the GW signal in a variety of ways, as detailed in Section V.5. Thus, using the spins and masses inferred from LIGO, Virgo and KAGRA measurements, the allowed boson mass can be constrained: if certain BH masses are detected with low spins, this could be evidence that boson clouds have spun down BHs. On the contrary, if BHs with high spins are detected, this could rule out the presence of boson clouds existing in a particular mass range Ng:2020ruv. Such an analysis allows us to obtain some evidence for or rule out the presence of boson clouds around rotating binary black holes, but does not necessarily permit us to claim confidently that we have detected a boson cloud.
Finally, unknown galactic BHs could take on a wide range of masses, ages and spins depending on when and how they formed. The number of expected BHs in the galaxy is $\mathcal{O}(10^{9})$, which comes from considering the age of the galaxy ($\sim 10^{11}$ years) and the supernova rate ($\sim 1$ per century) Madau:2014bja. Thus, the BHs we observe are such a small portion of the total number that exist in the local universe, and may not be representative of the full population. However, searches for such systems are entirely blind: we do not know any BH mass or spin parameters, nor do we know the location of such systems. While this may seem to be a daunting task, computationally efficient methods from the CW community have been leveraged to do these kinds of searches DAntonio:2018sff; Isi:2018pzk; Jones:2023fzz; Jones:2024fpg.
@@ VI.2 Search methods
VI.2.1 SD Excess power method
One approach to searching for boson clouds around rotating BHs is the excess-power method, which models the expected signal as approximately monochromatic over the entire observation time. Because the source could lie anywhere on the sky, the signal is modulated by the Earth’s motion, leading to a Doppler-shifted frequency. To recover such a signal, the data are transformed into time–frequency maps and corrected for different sky positions, so that a truly monochromatic signal would appear as a narrow line DAntonio:2018sff. This strategy is similar to that described in Section III.1.2, but here the modulation arises from the Doppler effect rather than from the random velocity distribution of individual DM particles. In practice, the analysis is repeated using different coherence times $T_{\text{FFT}}$, which control the width of the frequency bins. Shorter $T_{\text{FFT}}$ values are less sensitive overall but more robust to small frequency drifts, making the search less dependent on specific signal-model assumptions.
@@ VI.2.2 Viterbi
Another method that is more robust against stochastic frequency variations relies on Hidden Markov Models (HMMs) to find tracks in the time-frequency plane. The main idea is to model a time-frequency GW signal probabilistically as a Markov chain of transitions between “hidden” (unobservable) frequency states and use a detection statistic to relate “observed” frequency states with the hidden ones quinn2001estimation. In the framework for this method, the probability to jump to a new state at a given time only depends upon the previous hidden state. In total, there would be $N_{Q}^{N_{T}+1}$ possible paths through the hidden states, where $N_{Q}$ is the total number of hidden states and $N_{T}$ is the total number of times at which we have observed frequency states. The Viterbi algorithm provides a recursive, computationally efficient way of maximizing the probability that the hidden set of states is responsible for the observed sequence of states.
In Fig. 16, taken from Isi:2018pzk, we show the optimal path found by the Viterbi method for injected boson cloud signals in Gaussian noise that has an associated random walk of the frequency, which is meant to simulate unknown theoretical features of the signal. We can see that in both cases, with less and more variability of the signal, the Viterbi algorithm can find the appropriate track.
Recently, the Viterbi algorithm has also been adapted to search for GWs from annihilating vector boson clouds, which have timescales significantly shorter than in the scalar case Jones:2023fzz; Jones:2024fpg, and for which numerical waveforms must be used to gauge sensitivity Siemonsen:2019ebd; Siemonsen:2022yyf.
@@
Figure 16: Taken from Isi:2018pzk. This figure illustrates how the Viterbi HMM algorithm described in Section VI.2.2 works to identify tracks in the time-frequency plane that would arise from GW emission from annihilating boson clouds around rotating BHs. The physics behind this form of DM has been discussed in Section V.1. In particular, these plots show two examples of injected and recovered boson cloud signals with random walk frequency variations $|\delta f|\leq 0.1\Delta f$ (weak variation) and (b) $|\delta f|\leq 0.5\Delta f$ (strong variation), where $\Delta f=1/T_{\rm drift}$ is the frequency bin size for the coherence time of the analysis $T_{\rm drift}=8$ days. The top panels show the injected random walk frequency evolutions, which cannot be seen by eye in the time-frequency tracks. The Viterbi optimal path results in a good match to the injected signal path. Each step in time on the $x$-axis is 8 days. The signal has a frequency of $201.2$ Hz, a spin-up of $10^{-12}$ Hz/s, a randomly chosen sky position, and a strain amplitude of $5\times 10^{-26}$.
@@ VI.2.3 Cross-correlation and other methods
We have already discussed in Section III.1.1 cross correlating detector data to search for ultralight dark matter that could interact directly with GW interferometers; however, it is important to note that such a method could also be used to search for boson clouds around rotating BHs, both deterministically and for a stochastic GW background composed of the superposition of cosmological or astrophysical GW signals. Furthermore, virtually any method in the CW community Riles:2022wwz could be tuned to look for such boson cloud systems because such methods already analyze almost monochromatic signals focusing on spin-downs but could also extend their parameter spaces to include spin-ups.
@@ VI.3.1 All-sky search constraints
The BSD excess power method described in Section VI.2.1 has been used to analyze LIGO O3 data to search for the presence of annihilating scalar boson clouds around isolated BHs within our galaxy LIGOScientific:2021jlr; KAGRA:2022osp. This search was very computationally expensive, since each sky position had to be analyzed individually. No evidence for scalar boson cloud systems was found in either case, and thus upper limits were placed on the presence of boson clouds with different ages, spins, and distances away from earth in two different ways Palomba:2019vxe; Dergachev:2019wqa; LIGOScientific:2021jlr:
@@
Figure 17: Taken from LIGOScientific:2021jlr. Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, exclusion region of particular boson mass /BH mass pairs are derived assuming different ages and distances (left: 1kpc; right: 15 kpc) for BHs with initial spins of $\chi_{i}=0.9$. The physics behind this form of DM has been discussed in Section V.1. This is the most model-agnostic way to probe boson clouds around rotating BHs: the limits do not assume distributions over the mass, spin, distance of BHs in the Galaxy.
Constraints from the second way are shown in Fig. 18, in which a Kroupa BH mass distribution $f(m)\propto M_{\rm BH}^{-2.3}$ Elbert:2017sbr between $5,~50]M_{\odot}$ and $5,~100]M_{\odot}$ is assumed, as well as a uniform initial BH spin distribution between [0.2,0.9]. For a fixed BH age, the maximum distance reach as a function of boson mass indicates that at least 5% of GWs from BH/boson cloud systems would have been detected by the search, i.e. their strain amplitudes are larger than the minimum detectable amplitudes at a given confidence level (the upper limits) of the search (an example of this quantity is in Section III.1.2, Eq. 71. ). As expected, on average when $50M_{\odot}$ is the maximum BH mass considered, the distance that can be reached is also smaller, since the GW amplitudes scales strongly with the BH mass.
Figure 18: Taken from LIGOScientific:2021jlr. Using upper limits from the O3 LIGO search for GWs from annihilating boson cloud systems, the maximum distance that the search in LIGOScientific:2021jlr could reach at which at least 5% of a simulated BH population would have been detectable, that is, with a strain greater than the value of the upper limit, is shown here. The physics behind this form of DM has been discussed in Section V.1. To create this plot, Ref. LIGOScientific:2021jlr assumed a uniform spin distribution between [0.2,0.9], different ages, and a Kroupa mass distribution with maximum mass $50M_{\odot}$ (left) and $100M_{\odot}$ (right). These limits depend on the distributions assumed and are thus less agnostic than those presented in Fig. 17.
We note that the search constraints present in, for example, Steltner:2023cfk; Ming:2024dug; Covas:2024nzs, could also be cast in terms of all-sky or directed search constraints on scalar or vector boson clouds, but with particular caveats that would have be worked out depending on the actual search parameter space.
@@ VI.3.2 Cygnus X-1 constraints
Cygnus X-1 is an x-ray binary that is close by, has a large spin, $\sim 0.9$, and has relatively well-measured orbital parameters, which make it a reasonable BH to target. Since the sky position is known, searches based on the Viterbi algorithm discussed in Section VI.2.2 can use much longer coherence times than those used in LIGOScientific:2021jlr, thus permitting exquisite sensitivity towards almost monochromatic signals, while also being robust against theoretical uncertainties in the GW emission from Cygnus X-1, e.g. due to uncertainties on the measured semi-major axis, etc., due to the presence of an accretion disk, etc.
@@
Constraints on self-interacting DM using the search results on O2 data in Sun:2019mqb are shown only in the left-hand plot, since that search was designed to analyze 211 annihilation from Cygnus X-1. Furthermore, current and future ground-based interferometers will be able to place constraints on the axion decay constant at frequencies of $\mathcal{O}(100)$ Hz (left plot), while space-based interferometers, such as DECIGO and MAGIS-Space are required to probe transition signals at frequencies below 1 Hz (right panel). The figure also highlights the possibility of multi-band GW astronomy: it would be possible to detect a transition from the 322 to 211 energy levels in space-based observatories while simultaneously observing the CW arising from annihilation of bosons in the 211 energy level in ground-based interferometers Collaviti:2024mvh. Annihilation signals from the 322 level (middle plot) would be too weak to be detected by any of the considered current or future GW interferometers.
@@
Figure 19: Taken from Sun:2019mqb. Upper limits at 95% confidence on the strain amplitude of annihilating scalar boson clouds coming from the BH in Cygnus X-1. The physics behind this form of DM has been discussed in Section V.1. The black curve indicates the upper limits from the search on O2 data, with the red stars denoting upper limits derived in frequency bands with injections. The two different curves correspond to the signal amplitude with different choices for the (unknown) age of the system (lower curve: $10^{6}$ years; upper curve: $5\times 10^{6}$ years. The GW fine structure constant is colored, and the gray regions denote boson masses disfavored under the two different age assumptions. The parameters of Cygnux X-1 used to compute the amplitude curves are: $M=14.8M_{\odot}$, $\chi_{i}=0.99$, and $d=1.86$ kpc.
@@
Figure 20: Taken from Collaviti:2024mvh. Parameter space reachable by current and next-generation GW interferometers for (left) 211 annihilation, (middle) 322 annihilation, and (right) 322 $\rightarrow$ 211 transition signals from a scalar boson cloud around Cygnus X-1, assuming an age of $10^{5}$ yr and the parameters for Cygnus X-1 given in Fig. 19. The numbers 211, 322, and the 322 $\rightarrow$ 211 transition refer to specific energy levels of the cloud. The physics behind this form of DM has been discussed in Sections V.1 and V.2. The dashed curves divide the parameter space into three regimes: (1) top, above orange-dashed line: the strong self-interaction regime, (2) middle, between dashed-orange and dashed-blue lines: the moderate self-interaction regime, and (3) bottom, below the dashed-blue line: the gravitational (negligible self-interactions) regime. The white dotted line shows the expected relation between the axion and its decay constant. Darker colors here correspond to weaker GW signals primarily due to smaller occupation numbers in the energy states of the boson clouds. Constraints on 211 annihilation come from ground-based interferometers, while constraints from transitions will come from space-based interferometers. The 322 annihilation signal is too weak to be detected by any current our future observatory. The O2 constraints come from a real analysis of GW data and are thus representative of both the actual power spectral density of that dataset and an exclusion region on the plot.